# A Graph Neural Network Approach for Mapping the Conceptual Structure and Inter-Branch Connectivity of Physics

### v1

@@ 3.2 Computationally-Generated Cross-Domain Hypotheses
These findings demonstrate the framework’s capability to identify both established physical principles and novel mathematical analogies. The consistent identification of connections between Statistical Mechanics and other domains validates the framework’s ability to uncover the mathematical structures that unify disparate areas of physics.
3.3 Validation Through Negative Results and Methodological Controls
Equally informative as the model’s successful predictions are its “failures” and the spurious connections it learned to avoid. These negative results provide crucial validation, demonstrating that the framework is sensitive to genuine mathematical formalism rather than arbitrary correlations or philosophical links. We highlight three key types of such results.
  • Elimination of Spurious Links via Disambiguation: A critical validation of our methodology is demonstrated by the model’s behavior before and after semantic disambiguation. When trained on the raw, non-disambiguated graph, the model produced numerous high-confidence but physically nonsensical predictions. A prime example was a top-ranked link (Score >0.98>0.98) between AC Power (Electromagnetism) and the Adiabatic Process (Thermodynamics). The model correctly identified a strong mathematical link based on the shared variable ‘P’, but it was blind to the fact that ‘P’ represented Power in the first equation and Pressure in the second.
    After our semantic disambiguation protocol was applied—systematically renaming variables like ‘P’ to ‘P_power’ and ‘P_pressure’—this high-scoring false positive was entirely eliminated from the results. This demonstrates that our disambiguation step is not merely a preprocessing formality but a crucial component of the framework. It acts as an effective “auditor” of scientific notation, ensuring the final discoveries are grounded in genuine mathematical similarity rather than notational coincidence.
  • Absence of Purely Conceptual Links: The model correctly avoids predicting links between concepts that are deeply related physically but syntactically dissimilar. For instance, the connection between Maxwell’s equations and Special Relativity—a cornerstone of modern physics—received a low score. This is expected, as their standard formulations share almost no symbolic variables. This “failure” correctly demonstrates that the framework operates as a syntactic pattern recognizer, not a system with latent physical knowledge.
  • Low Scores for Distant Formalisms: Finally, the model correctly assigns low scores to pairs of equations that are conceptually linked but mathematically distinct. The connection between Newton’s Laws and the Schrödinger Equation, for example, scored low (0.621). This reflects not only their different mathematical structures but also aligns with our explicit exclusion of elementary mechanics from the final dataset to focus the model on the formalisms of modern physics. Similarly, the profound analogy between Thermodynamic and Information Entropy received only a moderate score (0.743), indicating the model cannot capture this deep connection without an explicit information-theoretic context.
These results, both the elimination of false positives and the presence of true negatives, validate that the model’s discoveries are grounded in the mathematical syntax of the provided equations, not in arbitrary correlations or a misunderstanding of the underlying physics.
@@ 4 Discussion
The results may support my hypothesis that a graph neural network, when combined with a methodologically sound graph representation, can identify statistically significant mathematical patterns in a corpus of physics equations. The model’s ability to achieve a robust and high AUC score ($0.9742\pm 0.0030$) after a semantic disambiguation and validation process indicates the discovered relationships may reflect true structural similarities in the mathematical formalism, not notational artifacts or model overfitting.
The discovered connections must be interpreted as syntactic analogies, not as direct physical or causal relationships. For instance, the link between Photon Energy (Quantum Mechanics) and the relativistic Four-Momentum invariant (Relativity) does not imply a new physical theory connecting them. Rather, it suggests a “meta-law”: a shared mathematical grammar used to describe critical phenomena, even when they occur in vastly different energy regimes.
The model is “semantically blind”— it recognizes syntax, not physical causality. Therefore, its predictions must be interpreted as hypotheses that require further theoretical investigation.
@@ 5 Conclusion and Future Work
I have developed and preliminarily validated a GNN-based framework capable of mapping the mathematical structure of physical law. Our proposed GAT model achieves state-of-the-art performance (AUC = 0.9742) on this novel task, demonstrably outperforming classical and established GNN baselines in a statistically-validated comparison. By addressing key methodological challenges—from semantic disambiguation to objective edge weight definition—we have shown that this approach can identify non-trivial, physically interesting, and statistically significant patterns.
@@ 5.1 Current Limitations and Future Directions
  • Database: The immediate next step is to expand the equation database to over 1,000 equations. This richer corpus will not only broaden the map but also enable the encoding of deeper mathematical structures—such as the presence of specific operators, derivatives, or symmetries—as features, moving the analysis from a syntactic to a more profound structural level.
  • Dynamic Methodological Refinements: Future work will explore more adaptive methodologies. This includes developing learnable edge-weighting schemes that go beyond a fixed linear combination and extending our disambiguation protocol to act as a general-purpose “auditor” for standardizing notational conventions across different scientific knowledge bases.
  • Advanced Statistical Analysis: Building on the rigorous FDR validation presented, future research can focus on more nuanced metrics that assess the novelty and non-triviality of discoveries. We also plan to explore methods from causal inference to investigate the potential directionality of the identified mathematical relationships.
  • Interactive Tools for Expert Validation: To bridge the gap between computational patterns and physical insight, we plan to develop interactive visualization tools that facilitate an expert-in-the-loop process to validate and interpret the most promising machine-generated analogies.
@@ 5.2 Broader Implications
This work opens several avenues for the broader scientific community, including applications in education, philosophy of science, and other scientific domains like chemistry and biology.
@@ 6 Data and Code Availability
All code, the cleaned dataset, and model weights will be available at: https://github.com/kingelanci/graphysics.
@@
Correspondence should be addressed to Massimiliano Romiti (massimiliano.romiti@acm.org).
@@ Acknowledgments
The author thanks the arXiv endorsers and the anonymous reviewers for their valuable feedback. Special gratitude to the ACM community for professional development resources.

### v2

@@ 3.3 Computationally-Generated Cross-Domain Hypotheses
These findings illustrate the framework’s capability to identify both mathematical analogues of established physical principles and novel syntactic analogies. The consistent identification of connections involving Statistical Mechanics as a central hub supports the hypothesis that the framework is learning a meaningful representation of the mathematical structures that connect different areas of physics.
4 Symbolic Analysis of Physics Equation Clusters: From Computational Patterns to Physical Insights
4.1 Overview and Significance
The core preliminary findings of this paper emerge from the symbolic analysis of 30 high-confidence equation clusters. This analysis reveals a hierarchy of insights, progressing from validating known physics to identifying errors and synthesizing complex principles. Figure 7 shows a typical dense cluster passed to this analysis stage.
Figure 7: An example of a high-density (0.615), high-core-number (6-core) conceptual cluster identified by the GNN, centered on relativistic principles. The framework isolates such structurally significant clusters for deeper symbolic analysis.
The analysis yielded 24 simplified expressions (80%), with the remaining 20% failing due to insufficient parseable equations or lack of valid substitutions. These computational results require expert interpretation to distinguish between physical insights and mathematical artifacts.
4.2 Foundational Results
4.2.1 The Klein-Gordon/Dirac Hierarchy (Cluster #5)
Cluster #5 connected the Klein-Gordon and Dirac equations. The system selected the Klein-Gordon equation as backbone:
$\left(\partial_{\mu}\partial^{\mu}+\frac{m^{2}c^{2}}{\hbar^{2}}\right)\psi=0$ (2)
where $\partial_{\mu}\partial^{\mu}$ is the d’Alembertian operator. The system performed the following substitutions from equations within the cluster:
  • $m=\gamma\hbar/c^{2}$ (from Dirac Equation - Relativistic Fermions)
  • $\psi=0$ (from Dirac Equation - Relativistic Fermions)
  • $\hbar=c^{2}m/\gamma$ (from Dirac Equation - Relativistic Fermions)
  • $c=(\gamma^{2}v^{r^{2}}/(\gamma^{2}-1))^{r^{-2}}$ (from Lorentz Factor)
The substitution $\psi=0$ reduces the expression to the identity True.
This result confirms the known relationship where the Dirac equation:
$(i\gamma^{\mu}\partial_{\mu}-mc/\hbar)\psi=0$ (3)
represents the “square root” of the Klein-Gordon equation. Applying the Dirac operator twice:
$(i\gamma^{\nu}\partial_{\nu}+m)(i\gamma^{\mu}\partial_{\mu}-m)\psi=(-\gamma^{\nu}\gamma^{\mu}\partial_{\nu}\partial_{\mu}-m^{2})\psi=0$ (4)
Using the anticommutation relations $\{\gamma^{\mu},\gamma^{\nu}\}=2\eta^{\mu\nu}$, this reduces to $(\partial_{\mu}\partial^{\mu}+m^{2})\psi=0$. Every Dirac solution satisfies Klein-Gordon (though not vice versa), a structure that historically led to the prediction of antimatter.
4.2.2 Maxwell’s Self-Consistency (Clusters #1, #2)
The two largest clusters (99 and 83 equations) centered on Maxwell’s equations, confirming the internal consistency of electromagnetic theory. The system identified complex relationships between the wave equation for electromagnetic fields and Maxwell’s equations through multiple substitutions, though the resulting expressions contain parsing artifacts that require further investigation. These clusters validate the framework’s ability to recognize the mathematical coherence of established physical theories.
4.2.3 Electromagnetic-Fluid Coupling (Cluster #8)
Cluster #8 revealed an unexpected synthesis between fluid dynamics and electromagnetism, identified by an earlier version of the code that still had parsing issues but nonetheless uncovered this intriguing connection (full report in Supplementary Materials). Starting with Reynolds number as backbone:
$Re=\frac{\rho vL}{\eta}$ (5)
The system applied the following substitutions from electromagnetic equations in the cluster:
  • E=Bv+F/qE=-Bv+F/q (from Lorentz Force (Complete))
  • $t=I\epsilon/L$ (from Inductance EMF)
  • $v=\epsilon/(BL)$ (from Motional EMF)
  • $t=-d\epsilon/d\Phi_{B}$ (from Faraday’s Law)
  • $t=-\epsilon/\Phi$ (from Lenz’s Law Direction)
This produced the simplified expression:
$\frac{\epsilon\rho}{B\eta}$ (6)
This result represents a dimensionless parameter coupling electromagnetic and fluid properties, analogous to the Magnetic Reynolds Number in magnetohydrodynamics, demonstrating how the framework can identify cross-domain mathematical structures even without understanding the underlying physics.
4.3 Error Detection as Knowledge Auditing
4.3.1 Dimensional Catastrophe (Cluster #4)
The four-momentum cluster exposed an error in the knowledge base. Starting with the relativistic invariant as backbone:
$\mu pp_{\mu}=c^{2}mr^{2}$ (7)
The system applied the following substitutions from equations in the cluster:
  • $p=E/c-cm^{2}$ (from Four-Momentum - Relativistic Invariant)
  • $m=\sqrt{E-cp}/c$ (from Four-Momentum - Relativistic Invariant)
  • $c=m_{e}$ (from Compton Scattering - Wavelength Shift)
This produced:
$\mu pp_{\mu}=m_{e}r^{2}\sqrt{E-m_{e}p}$ (8)
With residual: $m_{e}r^{2}\sqrt{E-m_{e}p}+\mu pp_{\mu}$
The critical error is the substitution $c=m_{e}$, which is dimensionally incorrect—equating velocity [L/T] with mass [M]. This error originates from a mis-parsed Compton scattering equation where the system incorrectly extracted a relationship between the speed of light and electron mass, likely from the Compton wavelength formula $\lambda_{C}=h/(m_{e}c)$ where $m_{e}$ and $c$ appear together but represent fundamentally different physical quantities.
4.3.2 Category Confusion (Cluster #16)
The framework combined Bragg’s Law for crystal diffraction:
$2d\sin\theta=n\lambda$ (9)
with substitutions from interference equations:
  • $m=d\sin\theta/\lambda$ (from Interference - Young’s Double Slit)
  • $a=m\lambda/\sin\theta$ (from Single Slit Diffraction (Minima))
  • $m=(d\sin\theta-\lambda/2)/\lambda$ (from Interference - Young’s Double Slit (dsi))
This produced the tautology True, but mixed different physical contexts—crystalline solids versus free space.
4.4 The Unexpected Value of Errors: Analog Gravity (Cluster #15)
Cluster #15 combined 11 equations from fluid dynamics and electromagnetism. Starting with Bernoulli’s equation:
$P+\frac{1}{2}\rho v^{2}+\rho gh=\text{constant}$ (10)
The system applied substitutions:
  • $h=v/(gr^{2})$ (from Torricelli’s Law)
  • $g=v/(hr^{2})$ (from Torricelli’s Law)
  • $c=(\mu_{r}-1)/\chi_{m}$ (from Magnetic Susceptibility)
  • $P=P_{0}+\rho gh$ (from Hydrostatic Pressure)
This produced:
$\frac{(\mu_{r}-1)^{2}}{h^{2}m^{2}}$ (11)
This mixes magnetic permeability with fluid dynamical variables—dimensionally inconsistent and physically incorrect.
However, this error points to analog gravity research, where:
  • Sound waves in fluids obey equations mathematically identical to quantum fields in curved spacetime
  • Fluid velocity fields create effective metrics analogous to gravitational geometry
  • Sonic horizons mirror black hole event horizons
  • Navier-Stokes equations in $(p+1)$ dimensions correspond to Einstein equations in $(p+2)$ dimensions
Near the horizon limit, Einstein’s equations reduce to Navier-Stokes, suggesting gravity might emerge from microscopic degrees of freedom.
4.5 Statistical Summary
Table 3: Cluster Analysis Results
Category Count Examples
Theory Validation 2 Klein-Gordon/Dirac, Maxwell consistency
Novel Synthesis 8 Transport phenomena, Reynolds-EMF
Tautologies 5 Planck’s Law circular substitutions
Dimensional Errors 2 $c=m_{e}$, knowledge base errors
Category Errors 3 Bragg/Young confusion
Provocative Failures 4 Analog gravity connection
Insufficient Data 6 No valid substitutions
Total 30 80% produced interpretable results
The framework operates at the syntactic level, recognizing mathematical patterns without understanding physical causality. This limitation, combined with parsing issues in early iterations, requires human interpretation to distinguish profound connections from tautologies or errors. However, even computational failures can suggest legitimate research directions, as demonstrated by the analog gravity connection.
@@ 5 Discussion
This proof of concept demonstrates a framework capable of systematic pattern discovery in physics equations. The framework functions as a computational lens revealing hidden structures and inconsistencies difficult to discern at human scale. It does not replace physicists, of course, but fully developed could augment their capabilities through systematic exploration of mathematical possibility space.
The framework’s value lies not in autonomous discovery but in its role as a computational companion—a tireless explorer of mathematical possibility space that surfaces patterns, errors, and unexpected connections for human interpretation. Even its failures are productive, transforming the vast combinatorial space of physics equations into a curated set of computational hypotheses worthy of expert attention.
The transformation of mathematical pattern into physical understanding remains fundamentally human. But by automating the pattern-finding, the framework frees physicists to focus on what they do best: interpreting meaning, recognizing significance, and building the conceptual bridges that transform equations into understanding of nature.
@@ 6 Conclusion and Future Work
In this preliminary work, I have developed and tested a GNN-based framework capable of mapping the mathematical structure of physical law, and acting as a computational auditor. My GAT model achieves high performance on this novel task, and the primary contribution of this work lies in the subsequent symbolic analysis. This analysis suggests the framework has a potential multi-layered capacity to: (i) verify the internal consistency of foundational theories, (ii) help debug knowledge bases by identifying errors and tautologies, (iii) synthesize mathematical structures analogous to complex physical principles, and (iv) provide creative provocations from its own systemic failures.
@@ 6.1 Current Limitations and Future Directions
This work, while promising, represents an initial step. The path forward is clear and focuses on several key areas:
  • Systematic Parameter Exploration: The framework requires systematic testing with different hyperparameter configurations to identify optimal settings for various physics domains. Different weight combinations in the importance scoring and clustering algorithms may reveal distinct classes of mathematical relationships, generating a vast number of hypotheses that require careful evaluation.
  • AI-Assisted Hypothesis Screening: The system currently generates hundreds of potential cross-domain connections, many of which are spurious or trivial. Future work should integrate large language models or other AI systems to perform initial screening of these computational hypotheses, filtering out obvious errors, tautologies, and dimensionally inconsistent results before human expert review. This would create a multi-stage pipeline: graph-based generation, AI screening, and expert validation.
  • Database Expansion: The immediate next step is to expand the equation database. A richer and broader corpus would enable the encoding of deeper mathematical structures, moving the analysis from a syntactic to a more profound structural level.
  • Generalization as a Scientific Auditor: Future work will focus on generalizing the framework beyond physics to other formal sciences. This includes refining the disambiguation protocol to act as a general-purpose “auditor” for standardizing notational conventions across different scientific knowledge bases.
  • Collaboration with Domain Experts: To bridge the gap between computational patterns and physical insight, future work must involve an expert-in-the-loop process. Collaboration with theoretical physicists is essential to validate, interpret, and build upon the most promising machine-generated analogies and audit reports.
@@ 6.2 Broader Implications
This work may open several avenues for the broader scientific community. As an auditing tool, it could potentially be used to systematically check the consistency of large-scale theories. As an educational tool, it might help students visualize the deep structural connections that unify different areas of science. More broadly, this research contributes to the emerging field of computational epistemology, developing methods to study the structure and coherence of scientific knowledge. Ultimately, this framework is presented as a tangible step toward a new synergy between human intuition and machine computation, where AI may serve as a tool to augment and stimulate the quest for scientific understanding.
@@ 7 Data and Code Availability
All code, cleaned dataset, and model weights available at: https://github.com/kingelanci/graphysics.
8 Supplementary Materials
Complete prediction distributions and analysis results available at the GitHub repository:
  • Full distribution of AUC ¿ 0.92 prediction scores
  • Complete symbolic analysis for all 30 clusters
  • Bootstrap validation logs
  • All generated cross-domain hypotheses (not just selected examples)
@@ 8 Supplementary Materials
Correspondence: Massimiliano Romiti (massimiliano.romiti@acm.org).
@@ Acknowledgments
The author thanks the arXiv endorsers for their support. Special gratitude to the ACM community for professional development resources and to online physics communities for discussions that helped clarify physical interpretations.
While AI tools were employed for auxiliary tasks such as code debugging, literature search, formatting assistance, and initial draft organization, all scientific content, analysis, interpretations, and conclusions are solely the work of the author. Any errors or misinterpretations remain the author’s responsibility.